OpenStudy (anonymous):

My answer is B Choose the point which is a solution to the system of inequalities: y _<_ -2x + 3 y < 1/2 x - 4 (2 points) Answer (4, -6) (-4, 6) (4, 6) (-4, -6)

Parth (parthkohli):

Is that \(y \le -2x + 3\)?

OpenStudy (anonymous):

yes:)

Parth (parthkohli):

Okay!

OpenStudy (anonymous):

thank you!!!!!!! :))))))))))))))))))

Parth (parthkohli):

A second, there is another inequality too.

OpenStudy (anonymous):

what?

Parth (parthkohli):

:( Your answer is actually incorrect.

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

-4, -6?

OpenStudy (anonymous):

would they both be negative?

Parth (parthkohli):

Is the second \(y \le \dfrac{1}{2}x - 4\)?

OpenStudy (anonymous):